# DREAMS: Preserving both Local and Global Structure in Dimensionality Reduction

### v1

@@ Acknowledgments
We thank Pierre Lambert and Cyril de Bodt for their help with the SQuadMDS package, Koshi Watanabe for making the StarMAP code available and Alex Diaz-Papkovich for providing the preprocessed 1000 Genomes Project data.
This work was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) via Germany’s Excellence Strategy (Excellence cluster 2064 “Machine Learning — New Perspectives for Science”, EXC 390727645), the Gemeinnützige Hertie-Stiftung, and the National Institutes of Health (UM1MH130981). The content is solely the responsibility of the authors and does not necessarily represent the official views of the National Institutes of Health.
Appendix A Appendix
A.1 Datasets
@@
For the MNIST dataset [Lecun et al., 1998], which was downloaded using the torchvision API, we used the first 50 principal components.
@@
The raw 1000 Genomes Project data [The 1000 Genomes Project Consortium, 2015] is available at https://ftp.1000genomes.ebi.ac.uk. This dataset contains 3 450 human genotypes. We got the preprocessed data from [Diaz-Papkovich et al., 2019] and [Diaz-Papkovich et al., 2023]; this is in an integer-valued data matrix with 53 999 features, containing values 0, 1, and 2, representing the number of alleles differing from a reference genome. We used PCA to reduce the number of features to 50.
A.2 DREAMS-CNE and DREAMS-CNE-Decoder
@@
Figure S1: Spearman correlation of pairwise distances (CPD, global metric) is plotted against kkNN recall (local metric). The figure shows the spectrum of DREAMS-CNE, using a regularizer with precomputed PCA embedding, and DREAMS-CNE-Decoder, using a regularizer with a linear decoder, across all seven datasets. The bigger scatter points display DREAMS-CNE and DREAMS-CNE-Decoder with their respective default regularization strengths (which achieved the highest average local-global score across all data sets) and reference methods. Here, InfoNC-$t$-SNE is used as the $t$-SNE backbone and corresponds to the regularization strength of $\lambda=0$ while PCA corresponds to the maximal regularization strength of $\lambda=1$. In panel g (MNIST) the scatter points of PCA and DREAMS-CNE-Decoder lie exactly on top of each other.
A.3 Visualizations of all datasets

### v2

@@ Acknowledgments
We thank Pierre Lambert and Cyril de Bodt for their help with SQuadMDS, Koshi Watanabe for making the StarMAP code available, and Alex Diaz-Papkovich for providing the preprocessed 1000 Genomes data.
This work was funded by the Gemeinnützige Hertie-Stiftung. DK is a member of the Germany’s Excellence cluster 2064 “Machine Learning — New Perspectives for Science” (EXC 390727645). The National Institutes of Health (UM1MH130981) funded initial phases of this work. The content is solely the responsibility of the authors and does not necessarily represent the official views of the National Institutes of Health.
Appendix
Appendix A Datasets
@@
The raw 1000 Genomes Project data [The 1000 Genomes Project Consortium, 2015] is available at https://ftp.1000genomes.ebi.ac.uk. This dataset contains 3 450 human genotypes. We got the preprocessed data from Diaz-Papkovich et al. [2019, 2023]; this is in an integer-valued data matrix with 53 999 features, containing values 0, 1, and 2, representing the number of alleles differing from a reference genome. We used PCA to reduce the number of features to 50.
We used 50 000 unprocessed random samples of the Mammoth dataset [Smithsonian Institution, 2020, Noichl, 2025] and likewise used the unprocessed Landsat Satellite [Srinivasan, 1993] dataset. For the Mammoth dataset, cluster labels (solely used for coloring the visualizations) were obtained by applying Agglomerative Clustering with 11 classes to the original data.
For the MNIST [Lecun et al., 1998], Fashion MNIST [Xiao et al., 2017], and CIFAR10 [Krizhevsky et al., 2009] datasets (downloaded using the torchvision API), we used the first 50 principal components. For Fashion MNIST, we additionally rescaled the original data to the interval $[0,1]$ before computing the first 50 principal components.
Appendix B Role of global embedding scaling in the DREAMS loss
In the regularization term
$\mathcal{R}(Y)=\frac{1}{n}\|Y-\alpha\widetilde{Y}\|_{F}^{2},$ (6)
the scaling parameter $\alpha$ is used to match the scale of the global reference embedding $\widetilde{Y}$ with that of $Y$. We considered two scaling strategies:
$\alpha_{n}=\frac{\|Y\|_{F}}{\|\widetilde{Y}\|_{F}},\quad\alpha_{p}=\frac{\langle Y,\widetilde{Y}\rangle_{F}}{\|\widetilde{Y}\|_{F}^{2}}.$ (7)
By the Cauchy–Schwarz inequality, $\alpha_{n}\geq\alpha_{p}$, with equality if and only if $Y$ is a positive rescaling of $\widetilde{Y}$. The first option, $\alpha_{n}$, rescales $\widetilde{Y}$ to match the scale of $Y$, preventing the regularization term from imposing an artificial scale onto the scale-sensitive $t$-SNE embedding. The second option, $\alpha_{p}$, is the $\alpha$ that minimizes the regularization term $\mathcal{R}(Y)$ for a given $Y$ and $\widetilde{Y}$ and is inspired by Procrustes analysis [Goodall, 1991, Gower, 1975].
Experimentally, we found that $\alpha_{p}$ performed slightly worse than $\alpha_{n}$ on the Tasic et al. dataset, leading to reduced local structure preservation as measured by a smaller KNN value (Table S1). While the regularization term was, by construction, smaller with $\alpha_{p}$, the KL divergence was slightly higher and the scale of the final embedding was further from $t$-SNE’s than with $\alpha_{n}$ (Figure S1). Consequently, we chose to use the scaling with $\alpha_{n}$ and conducted all subsequent experiments with this scaling method. Importantly, both scaling strategies worked much better than not using scaling at all and fixing $\alpha=1$ (Figure S1, Table S1).
Treating $\alpha$ as a constant during optimization yields the gradient
[…]
which pulls $Y$ towards the scaled global reference embedding $\alpha\widetilde{Y}$. In contrast, using $\alpha_{n}$ scaling and allowing the gradients to propagate through $\alpha$, yields the gradient
$\nabla_{Y}\mathcal{R}(Y)=\frac{2}{n}\bigg(Y-\alpha\tilde{Y}\bigg)-\frac{2}{n}\bigg(\frac{\langle Y,\tilde{Y}\rangle}{\|Y\|_{F}\|\tilde{Y}\|_{F}}-1\bigg)Y,$ (9)
with an additional second term that corresponds to a shrinkage of $Y$ towards the origin, since $\frac{\langle Y,\tilde{Y}\rangle}{\|Y\|_{F}\|\tilde{Y}\|_{F}}\leq 1$. As mentioned above, since the $t$-SNE loss is scale-sensitive, we do not want the regularizer to have a bearing on the scale of $Y$. Nevertheless, in the CNE variant of DREAMS, we propagated the gradient through $\alpha$ and still obtained a good trade-off between the InfoNC-$t$-SNE and PCA performance (Figure 4). We also verified that treating $\alpha$ as a constant during optimization in the CNE version did not lead to any noticeable difference in performance.
Figure S1: Regularization error, KL divergence, scaling parameter $\alpha$ and embedding scale of DREAMS during optimization using different or no scaling methods. Results are reported as the mean over four random seeds on the Tasic et al. dataset.
Table S1: Comparison of DREAMS performance across various alignment approaches for matching $\widetilde{Y}$ and $Y$ on the Tasic et al. dataset. We compare default DREAMS with DREAMS using a lower scaling factor $\alpha_{p}$, additionally translationally aligned the global reference $\tilde{Y}$ to $Y$ (treating the shift as constant), and full Procrustes analysis, which also contains a rotational alignment. The smaller scaling with $\alpha_{p}$ was detrimental, while translational and rotational alignment had no effect as the $t$-SNE loss is invariant to them.
Metric DREAMS ($\alpha_{n}$) $\alpha_{p}$ $\alpha_{p}$ + translation full Procrustes no scaling
KNN 0.37 0.35 0.35 0.35 0.27
CPD 0.89 0.89 0.89 0.89 0.53
Appendix C Runtime analysis
DREAMS’ adds a small overhead for computing the gradients of the regularizer compared to openTSNE. As a result, it runs bit slower than openTSNE and similarly fast as SQuadMDS-hybrid. The other methods were faster than DREAMS (Figure S2).
Figure S2: Runtime comparison. The bars indicate the mean runtime and standard deviation across four random seeds on the Tasic et al. dataset. For DREAMS and DREAMS-MDS, the reported times also include the runtime of the respective global reference embeddings (PCA and MDS).
Appendix D DREAMS-CNE and DREAMS-CNE-Decoder
@@
Figure S3: Spearman correlation of pairwise distances (CPD, global metric) is plotted against kkNN recall (local metric). The figure shows the spectrum of DREAMS-CNE, using a regularizer with precomputed PCA embedding, and DREAMS-CNE-Decoder, using a regularizer with a linear decoder, across all eleven datasets. The bigger scatter points display DREAMS-CNE and DREAMS-CNE-Decoder with their respective default regularization strengths (which achieved the highest average local-global score across all data sets) and local (InfoNC-$t$-SNE) and global (PCA) reference methods. Here, InfoNC-$t$-SNE is used as the $t$-SNE backbone and corresponds to the regularization strength of $\lambda=0$ while PCA corresponds to the maximal regularization strength of $\lambda=1$. In panel j (MNIST) the scatter points of PCA and DREAMS-CNE-Decoder lie exactly on top of each other.
Appendix E Intuition of local-global score
Figure S4: Spearman correlation of pairwise distances (CPD, global metric) is plotted against kNN recall (KNN, local metric) for different methods with marked local-global score spectrum. The figure shows the structure preservation results on the Tasic et al. and Packer et al. dataset.
Appendix F Sensitivity of kkNN recall to choice of $k$
Figure S5: Local structure preservation metric and combined local-global score for different numbers of neighbors $k$ in the kkNN recall. Displayed are the results on the Tasic et al. dataset. Except for SquadMDS-hybrid, the ranking of methods in terms of kkNN recall is insensitive to the exact value of $k$.
F.1 Visualizations of all datasets
@@
Figure S12: Visualizations of the Mammoth dataset [Smithsonian Institution, 2020, Noichl, 2025] with all considered methods. Colors represent cluster assignment of original data and are solely used for visual clarity but do not correspond to any ground truth labels.
Figure S13: Visualizations of the Landsat Satellite dataset [Srinivasan, 1993] with all considered methods.
Figure S14: Visualizations of the Fashion MNIST dataset [Xiao et al., 2017] with all considered methods.
@@
Figure S16: Visualizations of the CIFAR10 dataset [Krizhevsky et al., 2009] with all considered methods.